Short Report – QA/QC Meeting Apr 2018

Date: 2018-04-19

Participants / Excused

PM, EPL, PL, IB, CE, LH, KF, RM (8) / MG, TB, SB

Talking Points


  1. Please attend QA/QC meetings or tell LH if you can’t, needed for planning of the meetings.
  2. SRP & Tech-SRP: please send short statement about site to LH (max. 2 sentences)
  3. For digital data, (e.g. SWC @CH-OE2) should we store the unconverted raw data (float number) or the converted raw data (units, e.g. m3 m-3 for SWC) in our files? For now, the best solution is to store both variables (unconverted and converted raw data) in the raw data file. For later, we suggest to store only the unconverted raw data in a database. Then, we could easily convert it to units as needed. This conversion from unconverted to converted raw data would be a new step in the GL dataflow, i.e. it would be a new section in data handling, because the QuickViews are only for raw data as they arrive directly from the site (no conversions are done @ETH).
  4. Preview Swiss FluxNet homepage: online QA/QC checks, photo database


  1. [UPCOMING] Automatic fluxes will be activated again (goal is April 2018), priority conflicts w/ ICOS. 4 sites are now active.
  2. [UPCOMING] QA/QC Weekly checks in online checklists (instead of the server Excel sheet), starting w/ CH-DAV. May 2018.
  3. [UPCOMING] QV File Reports: checks available files on server. May 2018, if ICOS allows.

Site Reports


SRP (LH, fluxes): Still winter conditions, i.e. no consistent CO2 uptake yet. EC system seems to be running well, no gaps.
Tech-SRP (IB): –


SRP (KF): We observed increased uptake of CO2 during the last month. The first slurry application took place on the control parcel (Parzelle 4 in farmers jargon), but the QCL failed (and is now shipped to Aerodyne for troubleshooting) so it could not measure that event. 
Tech-SRP (PL): The fans of the TRH weather hut and CNR1 Net Radiometer had stopped working (for an unknown length of time) due to corroded electric contacts. The instruments were cleaned on the 18.04.18.


SRP (MG, LH): LI-7200 has still not returned from calibration. Site is currently made ready for ICOS (including the dataflow from the site to the ETC) and is therefore in a transition phase.
Tech-SRP (PM): LI-7200 and QCL not at site. No EC data. Everything else running smooth.


SRP (LH): EC and meteo systems are running fine, also no major gaps so far.
Tech-SRP (PL): The PR2 soil moisture profile probe is not measuring correct data (Werner compared the data to the existing Theta probe SWC profile). We are in contact with Delta-T to investigate the cause of the corrupted measurements. The instruments were cleaned on the 18.04.18.


SRP (EPL): Both CH-LAE and CH-LAS are running well. For CH-LAE, we can see a clear gradual increase in H2O fluxes starting in March. CO2 uptake from the trees not yet started (no leaves yet). For CH-LAS, nice diurnal CO2 uptake starting in March with wild garlic.
Tech-SRP (TB): (holidays)


SRP (CE): Suspicious sonic temperature patterns in the raw data still occur sporadically. Otherwise, everything is OK.
Tech-SRP (TB): (holidays)


You are here: Short Report - QA/QC Meeting Apr 2018


Last Updated on 30 May 2018 15:49

Leave a Reply

You must be logged in to post a comment.